Fermi National Laboratory
 

Accelerator Update for June 4, 2007 - June 6, 2007

 

Monday June 4

The day shift began with Operations monitoring Tevatron (TeV) store 5471 (57.81E30) and the Recycler stash (260.2E10), with the Antiproton Source (Pbar) stacking (26.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Interlock technicians placed a hurricane fan into Switchyard (SY) enclosure C at 8:24 AM to help deal with its ventilation and humidity problems. A Safety Officer lifted his access restriction from SY enclosure C for as long as the hurricane fan remained.

At 9:42 AM, a Linac quadrupole power supply (405) tripped off. An expert investigated.

Operations terminated store 5471 at 10:01 AM.

Operations prepared CDF and D0 for access at 10:24 AM.

A Pbar expert began an equalizer study that disrupted stacking around 10:29 AM.

Two MI experts began a slipstacking study at 10:46 AM.

At 10:47 AM, a Booster RF station (BRF12) went down. Operators investigated and discovered that a contract electrician working in the area had accidentally turned off the station. Operators brought the station back on line.

A new Pbar study began at 11:26 AM. Operations initiated a NuMI only time line in the Time Line Generator (TLG).

Pbar studies ended at 11:58 AM. They resumed normal stacking.

At 1:05 PM, a MI kicker tripped off. It reset.

The Neutron Therapy Facility (NTF) reported at 1:06 PM that they had completed treatment on their last patient for the day.

D0 came out of access at 1:48 PM.

At 2:08 PM, a CDF expert reported that they might have a cryo system leak on the detector platform in their collision hall. Water is coming from a known ice ball. They need to stay in access to investigate this further.

A MI expert began a 30-minute test of their MI flying wire system at 2:45 PM.

At 2:50 PM, a number of Linac Klystron and regular RF stations tripped off. Operations reported 258 Linac alarms. Later, a Linac expert reported that he had edited a file that was supposed to bypass device alarms; instead the file set the devices to bad values.

CDF came out of access at 3:16 PM.

Linac resumed sending beam at 3:34 PM.

Operations put the TeV through a dry squeeze at 3:53 PM.

Operations reported at 4:19 PM that the Recycler had stashed antiprotons once during the day shift.

Operations put the TeV into shot setup at 4:33 PM.

At 4:55 PM, Operations reported problems with MiniBooNE beam and the MI-8 collimators. Operators lowered the intensity of the MiniBooNE beam, which temporarily resolved the problem.

Operations established store 5473 at 7:05 PM with an initial luminosity of 227.1E30.

At 11:41 PM, KRF6 tripped off on a firing circuit problem. Operations contacted a Linac expert and then reset a VME crate. They brought KRF6 back on line at 11:52 PM.

Operations reported at 11:55 PM that the Recycler had stashed twice during the evening shift.

Tuesday June 5

The midnight shift began with Operations monitoring store 5473 (120.37E30) and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:06 AM, Operations reported that MIRF13 tripped off on a “Modulator Fan Off” indication and wouldn’t reset. Operators investigated, lowered the MI beam intensity, called an expert, but then the “Fan Off” indication cleared, and operators were able to turn the station back on and raise MI intensity back to its normal level.

At 1:21 AM, the NuMI beam permit tripped off due to bad vertical beam positions. It reset.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

Pbar experts began an equalizer study at 9:53 AM. Operations loaded a NuMI only time line into the TLG.

The Pbar equalizer study ended at 12:03 PM. Pbar resumed stacking.

MI/NuMI experts began a multibatch slipstacking study at 2:51 PM.

At 3:17 PM, the MI Power Supply (PS) Loop interlock tripped off. The Recycler lost 15E10 of the stash due to this trip. Experts investigated.

Operations reported at 3:59 PM that the Recycler had stashed two times during the day shift.

An MI expert reported at 4:24 PM that he replaced all the fuses for a repeater chassis and MI came back up.

The MI/NuMI slipstacking studies ended at 5:28 PM.

At 6:33 PM, a TeV RF station (T:RFST7) tripped off. Operations contacted an expert.

A TeV expert began a Tevatron Electron Lens (TEL1) study at 10 PM.

Operations terminated store 5473 at 11:59 PM.

Operations reported at midnight that the Recycler had stashed two times during the midnight shift.

Wednesday June 6

The midnight shift began with Operations monitoring the stash (341.4E10), with Pbar stacking (8.7E10), and with NuMI and MiniBooNE taking beam.

TeV RF experts reported at 12:47 AM that they had completed their work on TRF7.

At 12:56 AM, the MIPS loop tripped off and then reset itself. It was down for less than five minutes, but Recycler lost about 20E10 of its antiproton stash.

Operations put the TeV into shot setup at 1:03 AM.

At 1:33 AM, Operations reported that the P1 transfer line beam permit had tripped off twice due to a quadrupole power supply (Q703) going out of tolerance. When this power supply tripped off, it also caused MI and NuMI to trip off. Everything reset.

At 1:47 AM, a Pbar Debuncher RF station (DRF1) tripped off. It reset.

At 1:54 AM, BRF6 tripped off. It reset.

At 1:57 AM, BRF1 tripped off and would not reset. Operators bypassed the station and brought BRF13 on line.

At 3 AM, LRF4 tripped off. It reset.

At 3:05 AM, the TeV suffered a quench in sector C2 and A1. Operators investigated.

At 4:09 AM, the TeV sector A4 air compressor tripped off for the second time this shift. Operators reset the compressor, but it tripped off again. Operators and the Duty Mechanic tried greasing the compressor’s motor bearings, but it didn’t help. Operations contacted an expert. The TeV sector A4 beam valves closed due to low air pressure, but since the TeV had quenched, this wasn’t a problem.

Cryo system technicians at CHL reported at 4:21 AM that the TeV sector C2 cryo system cooldown was having trouble. The techs contacted the Cryo Coordinator. They reported at 4:54 AM that the C2 system was cooling down normally.

Operations reported at 5:17 AM that an expert was coming in to look at the TeV sector A4 air compressor.

Operations reported at 5:31 AM that a TeV expert was coming in to help investigate the cause of the quench.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to recover from the TeV quench and then resume stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/06/2007    email Fermilab